OpenStudy (anonymous):

A nuclear cooling tower cross section is a hyperbola having a diameter of 60 ft at the center. The distance between the two foci is 102 ft. What is the equation for the hyperbola?

OpenStudy (amistre64):

x^2 y^2 ---- - ---- = 1 900 b^2

OpenStudy (amistre64):

b^2 = 51^2 - 900

OpenStudy (amistre64):

b^2 = 1701